OpenStudy (anonymous):

find the eigen values and eigen vectors a=(-2 2 -3) (2 1 -6) (-1 -2 0)

ganeshie8 (ganeshie8):

did you solve the characteristic equation ?

OpenStudy (anonymous):

yes sir.........

ganeshie8 (ganeshie8):

good :) wat do you get for roots ?

ganeshie8 (ganeshie8):

\[ \left| A - \lambda I \right| = \left| \begin{array}{ccc} -2- \lambda & 2 & -3 \\ 2 & 1-\lambda & -6 \\ -1 & -2 & -\lambda \end{array} \right| = ?\]

ganeshie8 (ganeshie8):

if you have the roots (eigen values) for above, we can start from there :)

OpenStudy (anonymous):

-2 2 -3 2 1 -6 -1 -2 0

ganeshie8 (ganeshie8):

thats the given matrix, did u solve the characteristic equation yet ?